# Token Embeddings Violate the Manifold Hypothesis

### v2

@@ 3.2. Context does not persistently resolve singularities
Consequently, the contrapositive of Lemma 1 states that if $T$ is not a manifold but $T^{w-1}$ is a manifold, it follows that the space of sequences of tokens [^2] Caution: $T^{w}$ is the sequence of tokens embedded individually, which is different from the embedding of the entire sequence. Theorem 2 asserts conditions under which $T^{w}$ has a direct impact on LLM output. $T^{w}$ cannot be a manifold. At best, the latent space for every other context window size can be a manifold. For every proposed context window length, a longer window will exhibit singularities. In short, lengthening the context window does not persistently resolve singularities in the token embedding.
The remaining question is whether these singularities persist into the output of the LLM. Can the user elicit the instabilities caused by singularities? We provide sufficient conditions for unstable outputs:
Theorem 2.
Let $Z$ be the $d$-dimensional bounding manifold [^3] Since $T\subseteq\mathbb{R}^{\ell}$, it is always contained within such a bounding manifold $Z\subseteq\mathbb{R}^{\ell}$. for the token subspace, such that $T\subseteq Z$. Consider an LLM with a context window of size $w$, in which the latent space of tokens is $\mathbb{R}^{\ell}$, and we collect $m$ tokens as output from this LLM.
Suppose the following:
  1. (1)
    (enough tokens are collected from the response) $m>\frac{2wd}{\ell}$, but
  2. (2)
    (the context window is longer than the number of tokens we collected) $w\geq m$.
Under these conditions, a generic set of transformers yield a topological embedding of $T^{w}$ into the output of the LLM.
Proof.
See Section A.4
@@
3.3. Experimental results
@@
We applied our test to the token subspaces of the four open source LLMs introduced in Sec. 4.1: GPT2, Llemma7B, Mistral7B, and Pythia6.9B. In each LLM we tested, we found evidence that the token subspace is not globally a manifold because the local dimension is highly variable. Moreover, we can infer that they are either not fiber bundles or have high curvature locally; in either case—violating the hypotheses or having high curvature—model behavior near certain tokens will be highly unstable.
We observe highly statistically significant differences in the irregular tokens between LLMs—even for those with identical sets of tokens overall—which indicates that their respective training methodologies have a strong impact on the token subspace. Given these differences, prompts containing irregular tokens will likely produce dissimilar outputs across the four LLMs corroborating the results shown by [17]. Additionally, prompts containing irregular tokens within an LLM will likely be unstable and produce highly variable outputs with replicated queries. This is clear if $B(\psi)$ contains a singularity but will also be true for regions of high curvature resulting in instability for both interpolation and extrapolation tasks ([39]).
Several non-manifold features are visibly present in Figure 1. The tokens “ember” and “P” appear to be both putative examples of cusp points, places where the space exhibits a sharp exposed point. Manifolds cannot have cusp points, so attempting to approximate a cusp with a manifold—for instance, with UMAP—necessitates high curvature. The tokens “:” and “ly” appear to be pinch points—like what are present in a chain of sausages—which again cannot be approximated by a manifold unless it has high curvature.
Several additional views of Figure 1 are shown in Figure 6 in the Supplement. The coordinates, token text, $p$-values, and a video of the space with commentary are available in the ZIP file Supplement.
@@
Table 1. Dimensional data for and number of tokens rejecting the manifold and fiber bundle hypotheses at two slope changes. $n$ is the number of tokens (vocabulary size) in each model. The $p$-values displayed report the minimum $p$-value for each test, accounting for multiple testing. Quartiles (Q2 is the median) are shown for the distribution of the dimensions estimated from the slopes of the log-volume versus log-radius plots aggregated over all tokens. The “rejects” columns list the number of tokens rejecting each hypothesis. Table 3 (resp. Table 49) in the Supplementary material shows each of the fiber bundle (manifold hypothesis) violations for each model that is listed here.
Model Manifold Fiber bundle
rejects Smaller Radius Larger radius
@@
GPT2 $68$ Q1: $20$ $12$ Q1: $8$ $7$
$n=50257$ $p\approx 3\times 10^{-8}$ Q2: $389$ $p\approx 9\times 10^{-6}$ Q2: $14$ $p\approx 3\times 10^{-8}$
@@
Llemma7B $33$ Q1: $4096$ 00 Q1: $8$ $1$
$n=32016$ $p\approx 5\times 10^{-9}$ Q2: $4096$ N/A Q2: $11$ $p\approx 3\times 10^{-4}$
@@
Mistral7B $40$ Q1: $9$ $1$ Q1: $5$ $2$
$n=32016$ $p\approx 3\times 10^{-7}$ Q2: $48$ $p\approx 8\times 10^{-4}$ Q2: $6$ $p\approx 8\times 10^{-5}$
Q3: $220$ Q3: $9$
Pythia6.9B $54$ Q1: $2$ 00 Q1: $2$ 00
$n=50254$ $p\approx 2\times 10^{-7}$ Q2: $108$ N/A Q2: $5$ N/A
@@
Table 1 shows the results of our hypothesis tests for the four models we analyzed. It is clear that the models have quite different token input embeddings, and all of them exhibit highly significant rejections of the manifold hypothesis. GPT2, Llemma7B and Mistral7B also reject the fiber bundle hypothesis. The rejections of the fiber bundle hypothesis are more frequent in the larger radius region than the smaller radius region, which is consistent with the polysemy interpretation of [22].
While most of the tokens are not shared between the LLMs, Llemma7B and Mistral7B do have identical token sets. The fact that Table 1 shows significant differences between these two models indicates that the structure of the irregularities for these these two models is quite different. This implies that their response to the same prompt is expected to be markedly different, even without considering the differences in their respective transformer stages.
Since there are many rejections of the manifold hypothesis, we summarize some general trends.
  • The GPT2 tokens at irregularities are tokens that can only appear at the beginning of words.
  • The Pythia6.9B tokens at irregularities are nearly all word fragments or short sequences of text that are quite meaningless on their own.
  • The Llemma7B and Mistral7B tokens at irregularities are a combination of the previous two: either they can only appear at the beginning of words or they are word fragments.
Table 2. Testing the hypotheses of Theorem 2: Will singularities generically persist into the output? See statement of Theorem 2 for variable definitions. “Small” and “Large” refer to the small and large radius estimates of dimensions from Table 1.
@@
Estimates of the bounding manifold dimension used in Theorem 2 for the four LLMs we studied are in the “dimension” columns of Table 1, with the most stringent being the dimension for smaller radius. Table 2 shows the results of testing the hypotheses of Theorem 2 using the medians (Q2). Mistral7B, and Pythia6.9B satisfy the hypotheses of Theorem 2 using the small radius dimension estimates, but GPT2 and Llemma7B do not. All four LLMs satisfy the hypotheses of Theorem 2 if we instead use the larger radius dimension estimates. This means that it is likely that the transformers of Mistral7B, and Pythia6.9B are unable to resolve the singularities that we show exist in the token subspace, so singularities will persist in the output of the model. If the true bounding manifold for GPT2 or Llemma7B is smaller than the small radius estimate suggests, these too will likely produce singularities in their output.
4. Discussion
None of the four LLMs we studied are manifold globally given non-constant local dimension. Locally, the token subspace cannot be manifolds with large reach, and three of the four also cannot be fiber bundles with large reach. Irregularities—tokens that cause rejections of the manifold hypothesis—occur in different ways across all four LLMs. Additionally, irregularities corresponding to violations of the fiber bundle hypothesis are tokens whose neighborhoods exhibit a dependency between the large- and small-scale variability.
The grammatical role of tokens is likely a root cause for some of sensitivity of LLMs to their prompts that has been observed in the literature ([20, 21]), and may explain why “explaining LLM behavior” remains challenging and open. Most methods for explaining LLM behavior in terms of dynamical systems, for instance, derive their inferential power from assuming that the token subspace is a manifold ([40, 41, 42]). Our results indicate that these theoretical methods may not apply to current LLMs.
One might argue that context helps to disambiguate and “fix” these problems, but this intuitive argument seems not always possible as context alone is insufficient to resolve singularities. As we showed in Section 3.2, singularities in the token subspace persistently lead to singularities in the space of token sequences, and under appropriate conditions these propagate into the output of an LLM.
Irregularities—which may in fact be singularities—are noticeable in a PCA projection of the token subspace (Figure 1 and Figure 6). Intuitively, this means that the use of inner products to perform token de-embeddings by orthogonal projections may not be accomplishing the intended objectives. Even small perturbations within the latent space may result in instabilities in these projections, providing potential explanations for the instabilities in LLM outputs.
@@ 4. Discussion
In addition to token instability, the fact that token subspaces are not manifolds and may have small reach means that the geodesic distance between tokens can also be very unstable. While the distance along geodesics can be defined, it may not correlate with any sense of semantic distance between tokens. Furthermore, as [38] found, in most of the models, there are tokens with dimension 00 neighborhoods. These tokens are theoretically isolated, implying that the token subspace is disconnected and hence intrinsically infinitely far from other tokens.
@@
Singularities may arise either as artifacts of the training process or from features of the languages being represented. Consistent with the idea that polysemy may yield singularities ([22]), several of the tokens where the fiber bundle hypothesis is rejected are clear homonyms. For instance, both “affect” and “monitor” can be used as either nouns or verbs, and their meanings are different in these two roles. Tokens also form fragments of text, and so, a token may correspond to homonyms after the addition of a prefix or suffix. A token like “aunder” can be prefixed to yield the word “launder”, a word with multiple meanings of opposite sense. Specifically, one can “launder” clothing (positive connotation) or “launder” money (negative connotation). Several other tokens where the fiber bundle hypothesis is violated form words with substantially different meanings or grammatical roles upon adding a prefix or suffix. For instance, “wins” can function as a noun, as a verb, or form the adjective “winsome”.
@@
The differences in how the manifold and fiber bundle hypotheses are rejected across different LLMs suggest that the training methodology for each model leaves an indelible fingerprint (or gDNA coined by [17]). Making general assertions about LLMs without consideration of their training and evolution is likely fraught. Even between Llemma7B and Mistral7B, which have identical token sets, prompts likely cannot be “ported” from one LLM to another without significant modifications if they contain tokens near irregularities. This may have serious ramifications for interacting LLMs such as those powering multi-agentic frameworks being deployed today ([43, 44, 45]).
@@
A few clear patterns among tokens near irregularities are nevertheless noticeable. Tokens that begin a word or are a word fragment are often located at an irregularity. Additionally, in Llemma7B (but not Mistral7B) and Pythia6.9B the tokens with unusually low fiber dimension often contain non-printing or whitespace characters. This suggests that these models are quite sensitive to text layout, perhaps to the exclusion of more semantically salient features in the text. Given our findings, future experiments can be run to explore the impact of irregular tokens on the variability of responses produced by different LLMs.
@@
4.1. Limitations
While topological theory motivates our statistical test, only the slope changes that occur for radii less than the embedded reach correspond to rejections of the manifold or fiber bundle hypotheses. If the reach of a manifold is small, then it likely has a high local curvature, and measuring distances on that surface will be numerically unstable. As a result, the reach itself is difficult to estimate accurately from sampled data ([23, 46, 47]). Our implementation tests the radius below an estimated reach, determined by an inspection of the radius-vs-volume plots of a random sample of tokens. If the true reach is smaller than our estimates, the curvature of the manifolds that fit the data well, found by any method, will be high.
And lastly, we emphasize that these tests allow us to reject the structure assumed in our null hypothesis (smooth and regular), but do not permit us to infer an alternative. While rejecting regularity does imply singularity, it does not afford us any evidence as to the type of singularity. That must be determined post hoc if possible.
Acknowledgments

### v3

@@ 3.2 Context does not persistently resolve singularities
See Section A.5. ∎
@@
4 Methods
The tests we propose are implemented via Algorithm 1. The coordinates of the points (token embedding vectors, in Section 5.2) are fixed when we analyze the data, and they do not change from run to run. Algorithm 1 is fully deterministic, so there is no requirement of multiple runs. Centered at a given point, the sorted list of distances to all other points contains the radii at which the Monte Carlo estimate of the volume of a ball increases. The corresponding volumes are simply the indices of this list. Both tests estimate dimension as slope in the log-log progression of the entries versus the indices ($3$-point centered difference estimates via numpy.gradient()), and detect slope changes.
Algorithm 1 Manifold and fiber bundle tests
[…]
3:$W$: sliding window size
[…]
7:Set of dimension estimates
8:procedure ManifoldAndFiberBundleTest($x_{\bullet}$, $v_{min}$, $v_{max}$, $W$)
9:  Compute $n\times n$ pairwise distance matrix $D$ between all tokens
10:  for Each column of $D$ do $\triangleright$ Columns correspond to token indices
11:   Sort the column $\triangleright$ Now row indices of distance matrix are volumes, entries are radii
12:   Retain rows $v_{min}$ through $v_{max}$
13:   Compute log-log slopes (= dimension estimates) along the column
14:   Run two sample $T$-test along adjacent sliding windows of size $W$ with level $\alpha$:
Manifold test:
Append to $p_{1}$: the $p$-value for the hypothesis that the slope is constant
Fiber bundle test:
Append to $p_{2}$: the $p$-value for the hypothesis that the slope decreases with row index
15:   Store both $p$ values and slope with corresponding token (column index)
16:  end for
17:  Apply Holm-Bonferroni multiple test correction to both sets of $p$-values
18:end procedure
5 Results
To demonstrate the validity of our tests in practice, we demonstrate them on three synthetic datasets in Section 5.1 shown in Figure 3(a)–(c). We then turn to the main thrust of the paper, testing LLM token embeddings in Section 5.2. Source code is available at Robinson (2025).
5.1 Synthetic datasets
Figure 3: This figure shows the application of this paper’s manifold hypothesis test on three synthetic datasets: (a) a manifold, the unit sphere in $\mathbb{R}^{3}$, (b) a fiber bundle, the cusp surface $z^{8}+x^{2}+y^{2}=0$ in $\mathbb{R}^{3}$, (c) neither, a “lollipop” space, and (d) a neighborhood of the token ember in Mistral7B. Points are colored by the $p$ values of the manifold hypothesis test; darker colors mean reject the manifold hypothesis.
Figure 3(a) shows the result of applying our manifold hypothesis test to the unit sphere in $\mathbb{R}^{3}$. Since the sphere is a manifold, the $p$ value for every point tested is close to $1$, and no rejections of either hypothesis occur.
Figure 3(b) shows the result of applying the manifold test to a portion of the surface given by
$z^{8}+x^{2}+y^{2}=0,\text{ where }x^{2}+y^{2}\leq 1.$
This surface is a fiber bundle in which the base space is simply the origin, and the fiber spaces are the radial line segments tracing out from the origin to the edge of the disk. It is a manifold with boundary (not a manifold), and the origin $x=y=z=0$ is a cusp singularity. The number of points participating in the cusp and boundary is small ($41$ points) compared with the total number of points in the dataset ($1200$ points).
The manifold hypothesis test strongly detects the cusp singularity in Figure 3(b), with $p<5\times 10^{-8}$. The manifold test fails to detect the boundary ($p>0.001$). While using a larger $v_{max}$ in Algorithm 1 improves the test’s sensitivity to the point where the boundary is detected, it also reduces the specificity of the test for identifying which points are singularities. On the other hand, the fiber bundle test is not rejected at all, which is consistent with the space being a fiber bundle.
Finally, Figure 3(c) shows a sample of $3000$ points drawn from a space shaped like a candy “lollipop,” constructed by attaching a line segment (the “stick”) to a point on the surface of a hollow sphere (the “candy”). The lollipop is neither a manifold nor a fiber bundle. On the interior of the stick, the space is a $1$-dimensional manifold. The candy is a $2$-dimensional manifold, except where the stick attaches. The entire space has exactly two singular points. At the very end of the stick is a singular point (a boundary point). Where the stick attaches to the candy is another singularity; it is a transition from the $1$-dimensional stick to the $2$-dimensional candy.
Running the manifold hypothesis test on the lollipop reveals low $p$-values on points near the boundary of the stick ($75$ points) and at the transition ($58$ points) (both have $p<10^{-5}$). Due to random sampling, the exact points of singularity are not actually present in the dataset; these detections are near the two singularities. None of the points near the boundary point of the stick cause the fiber bundle hypothesis to be rejected with $p<10^{-5}$, while $2$ points near the transition cause the fiber bundle hypothesis to be rejected with $p<10^{-5}$. Since even one rejection is sufficient, the tests correctly identify that the lollipop is neither a manifold nor a fiber bundle.
5.2 LLM token embeddings
@@
We applied our Algorithm 1 to the token subspaces of the four open source LLMs introduced in Sec. 1.1: GPT2, Llemma7B, Mistral7B, and Pythia6.9B. In each LLM we tested, Table 2 exhibits evidence that the token subspace is not globally a manifold because the local dimension is highly variable (“dim.” columns). Moreover, we can infer that they are either not fiber bundles or have high curvature locally; in either case—violating the hypotheses or having high curvature—model behavior near certain tokens will be highly unstable (low $p$-values in the “rejects” columns). Tables 510 in the Supplementary material list every violation we found. Additionally, Figure 3(d) shows the application of the manifold test to the tokens of Mistral7B [^2] More views of Figure 3(d) are shown in Figure 8 in the Supplement. in the vicinity of the token ember. The manifold hypothesis is rejected at various tokens (low $p$-value, shown in darker color).
For each LLM we studied, we chose two pairs of $v_{min}$ and $v_{max}$ parameters in Algorithm 1: one yielding a small radius neighborhood and one yielding a larger radius neighborhood. These parameters were chosen by inspecting a small simple random sample of tokens, with the aim of identifying where changes in dimension were likely to occur. See Section A.6 for details. This configuration results in three tests: a single test for the manifold hypothesis, the fiber bundle test for small radius, and the fiber bundle test for large radius.
@@
Table 2: Dimensional data and number of tokens rejecting the manifold and fiber bundle hypotheses at two slope changes. $n$ is the number of tokens (vocabulary size) in each model. The $p$-values displayed report the minimum $p$-value for each test, accounting for multiple testing. Quartiles (Q2 is the median) are shown for the distribution of the dimensions estimated from the slopes of the log-volume versus log-radius plots aggregated over all tokens. The “rejects” columns list the number of tokens rejecting each hypothesis. Tables 410 in the Supplementary material show each of the manifold and fiber bundle hypothesis violations we found.
Model Manifold rejects Fiber bundle
Smaller Radius Larger radius
@@
GPT2 $n=50257$ $66$ $p\approx 3\times 10^{-8}$ Q1: 2020 $12$ $p\approx 9\times 10^{-6}$ Q1: 88 77 $p\approx 3\times 10^{-8}$
Q2: $389$ Q2: $14$
@@
Llemma7B $n=32016$ $33$ $p\approx 5\times 10^{-9}$ Q1: $4096$ 0 N/A Q1: 88 11 $p\approx 3\times 10^{-4}$
Q2: $4096$ Q2: $11$
@@
Mistral7B $n=32016$ $40$ $p\approx 3\times 10^{-7}$ Q1: 99 11 $p\approx 8\times 10^{-4}$ Q1: 55 22 $p\approx 8\times 10^{-5}$
Q2: $48$ Q2: 66
Q3: $220$ Q3: 99
Pythia6.9B $n=50254$ $54$ $p\approx 2\times 10^{-7}$ Q1: 22 0 N/A Q1: 22 0 N/A
Q2: $108$ Q2: 55
@@
Table 2 shows the results of our hypothesis tests for the four models we analyzed. The models consequently have token input embeddings with vastly different topology, and all of them exhibit highly significant rejections of the manifold hypothesis. GPT2, Llemma7B and Mistral7B also reject the fiber bundle hypothesis. The rejections of the fiber bundle hypothesis are more frequent in the larger radius region than the smaller radius region, which is consistent with the polysemy interpretation of Jakubowski et al. (2020).
The “rejects” columns of Table 2 give the number of tokens that fail our test (in the top rows of each cell). The listed $p$ is the overall $p$-value after Bonferroni correction. The “dim.” column shows the quartiles for the estimates of dimensions, at least for the tokens for which a dimension can reliably be ascribed. Note for each of the tokens listed in Tables 510, no dimension estimate is possible.
Since there are many rejections of the manifold hypothesis, we summarize some general trends, obtained by manual inspection of Tables 510 in the Supplementary material, to which we direct the reader. The GPT2 tokens at irregularities are tokens that can only appear at the beginning of words. The Pythia6.9B tokens at irregularities are nearly all word fragments or short sequences of text that are quite meaningless on their own. The Llemma7B and Mistral7B tokens at irregularities are a combination of the previous two: either they can only appear at the beginning of words or they are word fragments.
While most of the tokens are not shared between the LLMs, Llemma7B and Mistral7B have identical token sets. The fact that Table 2 shows significant differences between these two models indicates that the structure of the irregularities for these two models is quite different. This alone implies that their response to the same prompt is expected to be markedly different, even without considering the differences in their respective transformer stages, corroborating the results shown by Vargas et al. ((to appear).
Table 3: Testing the hypotheses of Theorem 2: Will singularities generically persist into the output? See statement of Theorem 2 for variable definitions. “Small” and “Large” refer to the small and large radius estimates of dimensions from Table 2.
@@
Estimates of the bounding manifold dimension used in Theorem 2 for the four LLMs we studied are in the “dimension” columns of Table 2, with the most stringent being the dimension for smaller radius. Table 3 shows the results of testing the hypotheses of Theorem 2 using the medians (Q2). Mistral7B and Pythia6.9B satisfy the hypotheses of Theorem 2 using the small radius dimension estimates, but GPT2 and Llemma7B do not. All four LLMs satisfy the hypotheses of Theorem 2 if we instead use the larger radius dimension estimates. This means that it is likely that the transformers of Mistral7B and Pythia6.9B are unable to resolve the singularities that we show exist in the token subspace, so singularities will persist in the output of the model. If the true bounding manifold for GPT2 or Llemma7B is smaller than the small radius estimate suggests, these too will likely produce singularities in their output.
6 Discussion
The large distribution of local dimension (Figure 1 and Table 2) rejects that the token embedding is a globally smooth manifold for each of the LLMs we studied. Beyond non-constant dimension, we also uncovered token neighborhoods without a well-defined intrinsic dimension, a structure not previously reported in these embeddings. These irregular structures suggest that the topology of the embeddings are more complicated than generally assumed. Consequently, all four token subspaces cannot locally be manifolds with large reach, and three cannot be fiber bundles with large reach. Given the aggregation of training data across domains (and even across natural languages), the potential for internal contradictions (contronyms) and confounders (natural language, tweets, source code), it is not surprising to see that our hypotheses are rejected.
The semantic role of tokens likely contributes to LLM sensitivity of prompts that has been observed in the literature (Schulz et al. (2025); Cui et al. (2025); Zhao et al. (2025a); Zilberman (2025); Sclar et al. (2023); Chao et al. (2023)) and may underscore why elucidating LLM behavior remains challenging and open. The manifold hypothesis gave both an intuitive and mathematically grounded explanation for the behaviors of learning machines (Liu et al. (2024); Brahma et al. (2015); Narayanan and Mitter (2010)) yet the instability of LLM outputs strongly suggested that these models were different. Our work gives conclusive empirical evidence against the manifold hypothesis for token embeddings and may offer novel ways to interpret LLM behavior as we are not in a regular setting.
@@ 6 Discussion
Such irregularities are noticeable in a principal components analysis (PCA) projection of the token subspace (Figures 3(d), 7 and 8) but are extremely hard to interpret a priori. Mathematicians are still classifying singularities in $\mathbb{R}^{3}$(Zhu et al. (2024)) implying that the classification problem is far from complete (even in $\mathbb{R}^{3}$). Intuitively, this means that the use of inner products to perform token de-embeddings by orthogonal projections may not be accomplishing the intended objectives. Even small perturbations within the latent space may result in instabilities in these projections, providing potential explanations for the instabilities in LLM outputs. In addition to token instability, the fact that token subspaces are not manifolds and may have small reach means that the geodesic distance between tokens can also be very unstable. While the distance along geodesics can be defined, it may not correlate with any sense of semantic distance between tokens. Furthermore, as Robinson et al. (2024) found, in most of the models, there are tokens with dimension 0 neighborhoods. These tokens are theoretically isolated, implying that the token subspace is disconnected and hence intrinsically infinitely far from other tokens.
@@
Singularities may arise either as artifacts of the training process or from features of the languages being represented. Consistent with the idea that polysemy may yield singularities (Jakubowski et al. (2020)), several of the tokens where the fiber bundle hypothesis is rejected are clear homonyms. For instance, both “affect” and “monitor” can be used as either nouns or verbs, and their meanings are different in these two roles. A token may also correspond to a pair of homonyms after the addition of a prefix or suffix. A token like “aunder” can be prefixed to yield the word “launder”, a word with multiple meanings of opposite sense. Specifically, one can “launder” clothing (positive connotation) or “launder” money (negative connotation). Several other tokens where the fiber bundle hypothesis is violated form words with substantially different meanings or grammatical roles upon adding a prefix or suffix. For instance, “wins” can function as a noun, as a verb, or form the adjective “winsome”.
@@
The differences in how the manifold and fiber bundle hypotheses are rejected across different LLMs suggest that the training methodology for each model leaves an indelible fingerprint (or gDNA coined by Vargas et al. ((to appear)). Making general assertions about LLMs without consideration of their training and evolution is likely fraught. Even between Llemma7B and Mistral7B, which have identical token sets, prompts likely cannot be “ported” from one LLM to another without significant modifications if they contain tokens near irregularities, likely a consequence of the invariance of domain theorem. This may have serious ramifications for interacting LLMs such as those powering multi-agentic frameworks being deployed today (Gu et al. (2025); Jin et al. (2025); Qian et al. (2025)).
@@
Tokens that begin a word or are a word fragment are often located at an irregularity. Additionally, in Llemma7B (but not Mistral7B) and Pythia6.9B, the tokens with unusually low fiber dimension often contain non-printing or whitespace characters. This suggests that these models are quite sensitive to text layout, perhaps to the exclusion of more semantically salient features in the text. Given our findings, future experiments can be run to explore the impact of irregular tokens on the variability of responses produced by different LLMs.
A natural question is if we can find a token embedding that is everywhere regular. One way to approach an answer is to view the number of tokens in a vocabulary as a hyperparameter. There might be a certain vocabulary size that minimizes the number of singularities given the corpus, architecture, etc. Even the best vocabulary might lead to a space that could still be singular, for reasons inspired by Jakubowski et al. (2020). Consider the * symbol, which happens to be singular for Mistral7B (See Table 8, number 29). In natural language, it has a very distinctive meaning in terms of calling out footnotes, while in other contexts it can be used for multiplication or marking text as bold (in Markdown). Therefore, this character connects these different semantics. Suppose that each language were in fact well-described by a manifold. Then these manifolds would be joined together in the entire latent space at the * (and likely other points). As a result, this token with high likelihood would be a singularity because the dimensions for different languages are almost surely different (Tulchinskii et al. (2023)). In reality, tokens are neither in 1:1 correspondence with words nor with grammatical symbols. Not only do parts of words map onto a single token, but tokenization is a many to one map across the board. How well context / attention / transformers can unfurl these mappings is a big part of the mystery of how these models work. Identifying and analyzing the irregularities may help us understand the approximate limits of attention/transformers architectures.
@@
7 Limitations
While topological theory motivates our statistical test, only the slope changes that occur for radii less than the embedded reach correspond to rejections of the manifold or fiber bundle hypotheses. If the reach of a manifold is small, then it likely has a high local curvature, and measuring distances on that surface will be numerically unstable. And lastly, we emphasize that these tests allow us to reject the structure assumed in our null hypothesis (smooth and regular), but do not permit us to infer an alternative. While rejecting regularity does imply singularity, it does not afford us any evidence as to the type of singularity. That must be determined post hoc if possible.
Acknowledgments and Funding Disclosure